# JLAB-THY-24-4244

### v1

@@ Abstract
We provide a detailed analysis of event-shape observables, namely the energy correlation function and jet angularity, for heavy-flavor jets produced in electron-positron collisions, focusing on quantum chromodynamics (QCD) interactions. Using modern jet substructure techniques, we investigate the dead-cone effect, where QCD radiation is suppressed around a heavy quark within an angle proportional to its mass. Our analysis achieves next-to-leading logarithmic accuracy, combined with fixed-order contributions, to improve the description of the transition near the dead-cone threshold. To ensure a comprehensive perspective, we compare our analytical results with predictions from the P YTHIA , H ERWIG , and S HERPA Monte Carlo simulations at past and future lepton colliders.
@@ I Introduction
In recent years, precision studies of jets with heavy flavors like charm ($c$) and bottom ($b$) quarks have gained momentum in the jet substructure community [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. However, at the moment, only a handful of measurements of jet substructure observables for jets seeded by heavy quarks is available [45, 46, 47, 45]. These considerations play a key role for Higgs boson studies and offer insights into the heavy-quark content of the proton by examining heavy-flavor jets recoiling against electroweak bosons. In particular, the recent development of flavor-jet algorithms [48, 49, 50, 51, 52] have opened the door to a novel yet-unexplored flavor-jet substructure program at high-energy colliders. From a theoretical view point, the mass of the heavy quarks sets a perturbative scale for the running coupling and at the same time acts as a regulator of the collinear divergences. Experimentally, the relatively long lifetimes of $B$- and $D$-hadrons cause their decay to occur away from the primary interaction point. Dedicated $b$- and $c$- tagging techniques that capitalize on this property to identify $B$ and $D$ hadrons or $b$ and $c$ jets are widely used in collider experiments [53, 54].
@@
To the best of our knowledge, only one study has measured jet substructure in $e^{+}e^{-}$ collisions [78]. These collisions offer the cleanest environment for jet production, free from multi-parton interactions (MPI), initial-state QCD radiation, and parton distribution function uncertainties. The absence of soft radiation from proton remnants uniquely allows the study of hadronization effects on jet substructure without interference from MPI -- unlike in $pp$ collisions, where these contributions overlap. Experimental investigations of $e^{+}e^{-}$ collisions can not only test the accuracy of theory predictions but also provide valuable data to refine non-perturbative models implemented in general-purpose Monte Carlo (MC) event generators such as P YTHIA  [79, 80, 81, 82, 83], H ERWIG  [84, 85, 86, 87], and S HERPA  [79, 80, 81, 88] as highlighted in [89, 90].
@@ III Fixed order analysis
To calculate the $\mathcal{O}(\alpha_{s})$ contribution $\Sigma_{V}^{(1)}$, we integrate Eq. (5) together with a $\Theta$ function that forces the observable to be smaller than a certain value $v$. In the resummation language, this corresponds to the first-order approximation of minus the radiator, $-\mathcal{R}^{(\text{f.o.})}_{V}$. We therefore evaluate
@@
×$\times\Big{[}\Theta(v-V(z,\theta^{2}))-1\Big{]},$ (7)
where $\xi=m^{2}/s$, and $k_{t}^{2}\simeq sz^{2}\theta^{2}/4$ with $\theta$ the angle between the emitting $b$-quark and the gluon. The upper limit $\theta_{\rm max}^{2}$ is irrelevant for our analysis as we focus on the logarithmic integration structure, on $\log\xi$ and $\log v$. The contribution $-1$ takes into account virtual corrections, while $V(z,\theta^{2})$ parameterises the observable in the quasi-collinear limit.
@@
¯$\bar{\mathcal{R}}^{(\text{f.o.})}_{V}(v,\xi)=$ $\frac{\alpha_{s}}{2\pi}\int^{\theta_{\rm max}^{2}}_{0}\frac{{\rm d}\theta^{2}}{\theta^{2}+4\xi}\int^{1-z_{\rm cut}}_{z_{\rm cut}}{\rm d}z\,P_{gb}\mathopen{}\mathclose{{}\left(z,k_{t}^{2}}\right)$
@@
¯$\bar{\mathcal{R}}^{\text{(f.o.)}}_{{\lambda}^{\alpha}}\mathopen{}\mathclose{{}\left(v,\xi}\right)$ $\simeq\mathcal{R}^{\text{(f.o.)}}_{{\lambda}^{\alpha}}\mathopen{}\mathclose{{}\left(v,\xi}\right)+\frac{\alpha_{s}C_{\text{F}}}{\pi}\Bigg{[}-\frac{1}{\alpha}\log^{2}\frac{\bar{v}}{z_{\rm cut}}\Bigg{]}.$ (10)
@@ IV All order analysis
[…]
@@
×$\times\frac{\alpha_{s}^{\text{CMW}}(\bar{k}_{t}^{2})}{2\pi}\Theta\mathopen{}\mathclose{{}\left(z\theta^{\alpha}-\bar{v}}\right),$ (11)
@@
×$\times\mathopen{}\mathclose{{}\left(1-\alpha_{s}\frac{\beta_{1}^{(4)}}{\beta_{0}^{(4)}}\frac{\log{\mathopen{}\mathclose{{}\left(1-\nu^{(4)}-\delta_{54}}\right)}}{1-\nu^{(4)}-\delta_{54}}}\right),$ (12b)
@@
×$\times\frac{\alpha_{s}^{\text{CMW}}(\bar{k}_{t}^{2})}{2\pi}\Theta\mathopen{}\mathclose{{}\left(z\theta^{\alpha}-\bar{v}}\right)\Theta(z-z_{\rm cut}).$ (14)
Both ungroomed and groomed radiators in Eqs. (IV) and (IV) are computed exploiting Lund diagrams for heavy flavors, introduced in [32]. We note that they are continuous in $\bar{v}=\xi^{\frac{\alpha}{2}}$, but their derivatives are not. This leads to discontinuous differential distributions at the transition. We can use fixed-order calculations to smooth this behavior and obtain a better theoretical description of the region near the dead-cone.
@@
×$\times\exp\mathopen{}\mathclose{{}\left[-2\mathopen{}\mathclose{{}\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,\xi)-R_{b}^{\text{(f.c.)}}(v,\xi)}\right)}\right],$ (15)
@@
×$\times\exp\mathopen{}\mathclose{{}\left[-2\mathopen{}\mathclose{{}\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,\xi)-R_{b}^{\text{(f.c.)}}(v,\xi)}\right)}\right].$ (16)
@@ V Numerical predictions
Figure 1: Ratio $\Sigma^{b}/\Sigma^{q}$ at parton level at LEP center-of-mass collision energy. Our NLL predictions are in black, while P
@@
MC predictions are in color. Vertical dashed black line marks the dead-cone transition threshold $\theta_{D}=2m_{b}/\sqrt{s}\approx 0.11$.
We now compare our NLL predictions with parton-level MC simulations from P YTHIA , H ERWIG , and S HERPA . The results are shown in Fig. 1 with solid red, green, and blue lines. For ungroomed observables, MC predictions differ significantly due to varying implementations of heavy particle radiation, see Refs. [110, 104], [111, 112, 113], and [114, 115], respectively. Notably, each MC overestimates the dead-cone effect, as $\Sigma^{b}/\Sigma^{q}$ increases before the dead-cone boundary.
The predictions for the groomed observables are given in the lower half of Fig. 1. We used mMDT grooming with $z_{\text{cut}}=0.1$. The NLL results for groomed $\Sigma^{b}/\Sigma^{q}$ ratios for $v<\theta_{D}$ show smaller enhancements compared to ungroomed ones. The grooming procedure also changes the slope and reduces discrepancies around the dead-cone threshold between NLL and MC predictions. Our NLL predictions for ECF and jet angularities differ at $v=\theta_{D}$. The $\Sigma^{b}/\Sigma^{q}$ for ECF is almost equal to unity, while jet angularities show a non-negligible deviation, making jet angularity a better observable for heavy-flavor studies.
Figure 2: Ratio $\Sigma^{b}/\Sigma^{q}$ at parton (solid lines) and hadron (dashed lines) levels at LEP center-of-mass collision energy simulated with P
@@
MC. Vertical dashed black line marks the dead-cone transition threshold $\theta_{D}=2m_{b}/\sqrt{s}\approx 0.11$.
@@
To estimate the impact of non-perturbative corrections on our results, we compare parton-level MC simulations (solid lines) with MC simulations that include hadronization corrections (dashed lines) in Fig. 2. As noted in Ref. [27], the jet substructure of $b$-jets is influenced not only by non-perturbative corrections but also by $B$-hadron decays. In agreement with predictions of Ref. [27] our previous $pp$ study [37] demonstrated that decay products of $B$-hadrons can significantly alter jet shapes. While such effects could be incorporated into our analysis using transfer matrices, as introduced in Ref. [90], we opt to simplify this study by treating all $B$-hadrons as stable in P YTHIA , H ERWIG , and S HERPA simulations. As a result, our findings can be directly compared to measurements where $B$-hadrons are reconstructed from their decay products.
@@
Let us delve into the results shown in Fig. 2. For the ungroomed case, hadronization significantly impacts the ratios $\Sigma^{b}/\Sigma^{q}$. Notably, the solid (parton-level) and dashed (hadron-level) curves for all three MC generators -- P YTHIA , H ERWIG , and S HERPA -- begin to diverge well before reaching the dead-cone threshold. Furthermore, these generators exhibit distinct predictions not only at the parton level (expected due to differences in their parton shower models) but also at the hadron level. The hadron-level discrepancies are usually smaller, as these MC generators have been tuned to provide consistent descriptions of existing data. This divergence highlights potential areas for improvement in current MC models and underscores the value of analyzing archived LEP data to refine MC tunes further.
@@
Now, let us focus on the groomed ratios shown in the lower half of Fig. 2. The application of the mMDT groomer significantly reduces the discrepancies between parton-level and hadron-level results, particularly near the dead-cone threshold. Notably, the ECF observable exhibits reduced sensitivity to non-perturbative corrections around the dead-cone boundary compared to the jet angularity. This highlights the effectiveness of grooming techniques in mitigating non-perturbative effects and improving the robustness of the analysis.
@@
To conclude this section, let us estimate the feasibility of observing the dead cone effect with ECF and jet angularities at future $e^{+}\,e^{-}$ colliders. In order to do that in Fig. 3 we compare our resummed predictions versus MC simulations at $\sqrt{s}=2$ TeV center-of-mass collision energy which corresponds to the ILC setup. The increase in the collision energy from 91 GeV to 2 TeV implies that for the MC results the ratios $\Sigma^{b}/\Sigma^{q}$ start to deviate from unity at much smaller values, roughly at $v\sim 0.1$. Similarly, the deadcone-driven changes become visible if $v\leq 0.005$ which is in agreement with the leading-log estimate of the dead-cone boundary. Such small thresholds might present a significant challenge for future ILC measurements, as detecting the dead cone effect with jets seeded by $b$-quarks would require exceptional detector resolution. Thus, we propose that analyzing archived LEP data may offer greater value for studying $b$-quark jet substructure than relying solely on future high-energy experiments.

### v2

@@ Abstract
We provide a detailed analysis of event-shape observables, namely the energy correlation function and jet angularity, for heavy-flavor jets produced in electron-positron collisions, focusing on quantum chromodynamics (QCD) interactions. Using modern jet substructure techniques, we investigate the dead-cone effect, where QCD radiation is suppressed around a heavy quark within an angle proportional to its mass. Our analysis achieves next-to-leading logarithmic accuracy, combined with partial fixed-order contributions, to improve the description of the transition near the dead-cone threshold. To ensure a comprehensive perspective, we compare our analytical results with predictions from the P YTHIA , H ERWIG , and S HERPA Monte Carlo simulations at past and future lepton colliders.
@@ I Introduction
In recent years, precision studies of jets with heavy flavors like charm ($c$) and bottom ($b$) quarks have gained momentum in the jet substructure community [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. However, at the moment, only a handful of measurements of jet substructure observables for jets seeded by heavy quarks is available [42, 43, 44, 42]. These considerations play a key role for Higgs boson studies and offer insights into the heavy-quark content of the proton by examining heavy-flavor jets recoiling against electroweak bosons. In particular, the recent development of flavor-jet algorithms [45, 46, 47, 48, 49] have opened the door to a novel yet-unexplored flavor-jet substructure program at high-energy colliders. From a theoretical view point, the mass of the heavy quarks sets a perturbative scale for the running coupling and at the same time acts as a regulator of the collinear divergences. Experimentally, the relatively long lifetimes of $B$- and $D$-hadrons cause their decay to occur away from the primary interaction point. Dedicated $b$- and $c$- tagging techniques that capitalize on this property to identify $B$ and $D$ hadrons or $b$ and $c$ jets are widely used in collider experiments [50, 51].
@@
To the best of our knowledge, only one study has measured jet substructure in $e^{+}e^{-}$ collisions [75]. These collisions offer the cleanest environment for jet production, free from multi-parton interactions (MPI), initial-state QCD radiation, and parton distribution function uncertainties. The absence of soft radiation from proton remnants uniquely allows the study of hadronization effects on jet substructure without interference from MPI - unlike in $pp$ collisions, where these contributions overlap. Experimental investigations of $e^{+}e^{-}$ collisions can not only test the accuracy of theory predictions but also provide valuable data to refine non-perturbative models implemented in general-purpose Monte Carlo (MC) event generators such as P YTHIA  [76, 77, 78, 79, 80], H ERWIG  [81, 82, 83, 84], and S HERPA  [85] as highlighted in [86, 87].
@@ III Fixed order analysis
To calculate the $\mathcal{O}(\alpha_{s})$ contribution $\Sigma_{V}^{(1)}$, we integrate Eq. (5) together with a $\Theta$ function that forces the observable to be smaller than a certain value $v$. In the resummation language, this corresponds to the first-order approximation of minus the radiator, $\mathcal{R}^{(\text{f.o.})}_{V}.$ We therefore evaluate
@@
$\times\Big{[}\Theta(v-V(z,\theta^{2}))-1\Big{]},$ (7)
where $\xi=m^{2}/s$, and $k_{t}^{2}\simeq sz^{2}\theta^{2}/4$ with $\theta$ the angle between the emitting $b$-quark and the gluon. The upper limit $\theta_{\rm max}^{2}$ is irrelevant for our analysis as we focus on the logarithmic integration structure, on $\log\xi$ and $\log v$. The contribution 1-1 takes into account virtual corrections, while $V(z,\theta^{2})$ parameterises the observable in the quasi-collinear limit.
@@
$\bar{\mathcal{R}}^{(\text{f.o.})}_{V}(v,\xi)=$ $\frac{\alpha_{s}}{2\pi}\int^{\theta_{\rm max}^{2}}_{0}\frac{{\rm d}\theta^{2}}{\theta^{2}+4\xi}\int^{1-z_{\rm cut}}_{z_{\rm cut}}{\rm d}z\,P_{gb}\mathopen{}\mathclose{{\left(z,k_{t}^{2}}}\right)$
@@
$\bar{\mathcal{R}}^{\text{(f.o.)}}_{{\lambda}^{\alpha}}\mathopen{}\mathclose{{\left(v,\xi}}\right)$ $\simeq\mathcal{R}^{\text{(f.o.)}}_{{\lambda}^{\alpha}}\mathopen{}\mathclose{{\left(v,\xi}}\right)+\frac{\alpha_{s}C_{\text{F}}}{\pi}\Bigg{[}-\frac{1}{\alpha}\log^{2}\frac{\bar{v}}{z_{\rm cut}}\Bigg{]}.$ (10)
@@ IV All order analysis
We now present the calculation to all perturbative orders, which enables the resummation of large logarithmic contributions due to the various scale hierarchies. We start with ungroomed jets and then move on to the mMDT groomed jets. We present our calculation that resums the logarithms of $v$ and the logarithms of $\xi$ in each of the regions defined by the hierarchy between $v$ and $\xi$ with NLL accuracy. This leads to a simplification and we only need to calculate one radiator, i.e. $\mathcal{R}_{V}(v,\xi)\xrightarrow{\text{NLL}}R(v,\xi)$.
@@
$\times\frac{\alpha_{s}^{\text{CMW}}(\bar{k}_{t}^{2})}{2\pi}\Theta\mathopen{}\mathclose{{\left(z\theta^{\alpha}-\bar{v}}}\right),$ (11)
@@
$\times\mathopen{}\mathclose{{\left(1-\alpha_{s}\frac{\beta_{1}^{(4)}}{\beta_{0}^{(4)}}\frac{\log{\mathopen{}\mathclose{{\left(1-\nu^{(4)}-\delta_{54}}}\right)}}{1-\nu^{(4)}-\delta_{54}}}}\right),$ (12b)
@@
$\times\frac{\alpha_{s}^{\text{CMW}}(\bar{k}_{t}^{2})}{2\pi}\Theta\mathopen{}\mathclose{{\left(z\theta^{\alpha}-\bar{v}}}\right)\Theta(z-z_{\rm cut}).$ (14)
Both ungroomed and groomed radiators in Eqs. (IV) and (IV) are computed exploiting Lund diagrams for heavy flavors, introduced in [32]. We note that they are continuous in $\bar{v}=\xi^{\frac{\alpha}{2}}$, but their derivatives are not. This leads to discontinuous differential distributions at the transition. We can employ fixed-order calculations, particularly the non-logarithmic contributions from Eqs. (III) and (III), to smooth this behavior and obtain a better theoretical description of the region near the dead-cone.
@@
$\times\exp\mathopen{}\mathclose{{\left[-2\mathopen{}\mathclose{{\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,\xi)-R_{b}^{\text{(f.c.)}}(v,\xi)}}\right)}}\right],$ (15)
@@
$\times\exp\mathopen{}\mathclose{{\left[-2\mathopen{}\mathclose{{\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,\xi)-R_{b}^{\text{(f.c.)}}(v,\xi)}}\right)}}\right].$ (16)
@@ V Numerical predictions
All resummed contributions indicated by solid black lines are supplemented with an uncertainty band. This band is calculated by rescaling the argument of the strong coupling by $x_{R}$ and rescaling the observable by $x_{v}$. Appropriate counter terms are added to maintain NLL accuracy. To estimate the theoretical uncertainty in the region around the dead-cone transition, we vary both $x_{R}$ and $x_{v}$ within specific ranges. We consider the cases where $x_{R}=1$ and $\frac{1}{2}\leq x_{v}\leq 2$, and where $\frac{1}{2}\leq x_{R}\leq 2$ and $x_{v}=1$. We obtain the uncertainty band by choosing the lowest and highest envelope out of five possible combinations, with the central curve being $(x_{R}=1,x_{v}=1)$.
@@
To demonstrate the accuracy of our collinear approximation, we also plot the dashed black line, which represents the truncation of the ratio $\Sigma^{b}/\Sigma^{q}$ at $\mathcal{O}(\alpha_{s})$.-The yellow line, on the other hand, represents the full real radiation matrix element at next-to-leading order (NLO) generated using M ADGRAPH 5_aMC@NLO [107]. For values of the observable $v\gtrsim\theta_{D}$, we find that our results approximate the full matrix element with an uncertainty of up to 10-20%.
Figure 1: The ratio $\Sigma^{b}/\Sigma^{q}$ at the parton level at LEP center-of-mass collision energy is presented. Our NLL predictions are in solid black, with their truncation at $\mathcal{O}(\alpha_{s})$ indicated by dashed black lines and the full real matrix element at NLO in yellow obtained using M
ADGRAPH
5_aMC@NLO. The predictions from P
@@
MC simulations are in color. The vertical dashed black line marks the dead-cone transition threshold $\theta_{D}=2m_{b}/\sqrt{s}\approx 0.11$. All MC simulations are performed at LO+PS accuracy.
We now compare our NLL predictions with parton-level MC simulations from P YTHIA , H ERWIG , and S HERPA . We note that all our MC simulations are performed using their default settings, and at LO+PS accuracy. The results are shown in Fig. 1 with solid red, green, and blue lines. For ungroomed observables, MC predictions differ significantly due to varying implementations of heavy particle radiation, see Refs. [108, 101], [109, 110, 111], and [112, 113], respectively. Notably, each MC overestimates the dead-cone effect, as $\Sigma^{b}/\Sigma^{q}$ increases before the dead cone boundary.
The predictions for the groomed observables are given in the lower half of Fig. 1. We used mMDT grooming with $z_{\text{cut}}=0.1$. The NLL results for groomed $\Sigma^{b}/\Sigma^{q}$ ratios for $v<\theta_{D}$ show smaller enhancements compared to ungroomed ones. The grooming procedure also changes the slope and reduces discrepancies around the dead-cone threshold between NLL and MC predictions. Our NLL predictions for ECF and jet angularities differ at $v=\theta_{D}$. The $\Sigma^{b}/\Sigma^{q}$ for ECF is almost equal to unity, while jet angularities show a non-negligible deviation.
Figure 2: The ratio $\Sigma^{b}/\Sigma^{q}$ at parton (solid lines) and hadron (dashed lines) levels at LEP center-of-mass collision energy simulated with P
@@
MC is presented. The vertical dashed black line marks the dead-cone transition threshold $\theta_{D}=2m_{b}/\sqrt{s}\approx 0.11$. The blue bands estimate the uncertainty due to two different hadronization models implemented in S
HERPA
To estimate the impact of non-perturbative corrections on our results, we compare parton-level MC simulations (solid lines) with MC simulations that include hadronization corrections (dashed lines) in Fig. 2. Currently, there are two primary models of hadronization available on the market: the Lund string model [79], as implemented in P YTHIA , and the cluster model, as implemented in H ERWIG  [83]. Unlike P YTHIA and H ERWIG , S HERPA incorporates both models with some modifications, as detailed in Ref. [85]. Consequently, by altering the hadronization models in S HERPA , one can estimate the overall uncertainty associated with hadronization effects. This is illustrated in Fig. 2, where the blue bands represent the uncertainty, with upper edge corresponding to the Lund string model and the lower one to the cluster model [^1] The rope hadronization model, developed by the P YTHIA collaboration [114], in principle allows us to perform a similar estimate of hadronization uncertainty using P YTHIA . However, we note that this model is relatively new and has not yet been tuned to the data. Therefore, we refrain from using it in the current study..
@@
Also, as noted in Ref. [27], the jet substructure of $b$-jets is influenced not only by non-perturbative corrections but also by $B$-hadron decays. In agreement with predictions of Ref. [27] our previous $pp$ study [37] demonstrated that decay products of $B$-hadrons can significantly alter jet shapes. While such effects could be incorporated into our analysis using transfer matrices, as introduced in Ref. [87], we opt to simplify this study by treating all $B$-hadrons as stable in P YTHIA , H ERWIG , and S HERPA simulations. As a result, our findings can be directly compared to measurements where $B$-hadrons are reconstructed from their decay products.
@@
Let us delve into the results shown in Fig. 2. For the ungroomed case, hadronization significantly impacts the ratios $\Sigma^{b}/\Sigma^{q}$. Notably, the solid (parton-level) and dashed (hadron-level) curves for all three MC generators - P YTHIA , H ERWIG , and S HERPA - begin to diverge well before reaching the dead-cone threshold. Furthermore, these generators exhibit distinct predictions not only at the parton level (expected due to differences in their parton shower models) but also at the hadron level. The hadron-level discrepancies are usually smaller, as these MC generators have been tuned to provide consistent descriptions of existing data. However, as evident in the figure, the H ERWIG yields predictions different from P YTHIA and S HERPA . Moreover, the S HERPA hadron-level uncertainty band does not overlap with its parton-level predictions. Consequently, we conclude that in the case of the ungroomed observable, the divergences between different event generators can highlight potential areas for improvement in current MC models and underscore the significance of analyzing archived LEP data to refine MC tunes further. It’s worth noting that such a pronounced influence of non-perturbative physics may effectively screen the essentially perturbative dead-cone effect for the setup considered in this paper.
@@
Now, let’s focus on the groomed ratios presented in the lower half of Fig. 2. The application of the mMDT groomer substantially reduces the discrepancies between parton-level and hadron-level results, particularly near the dead-cone threshold. Furthermore, the S HERPA hadronization uncertainty bands now overlap with parton-level S HERPA predictions, as well as with those of P YTHIA and H ERWIG . Notably, this effect is particularly pronounced for the case of the ECF observable, which exhibits reduced sensitivity to non-perturbative corrections around the dead-cone boundary compared to the jet angularity. This underscores the efficacy of grooming techniques in mitigating non-perturbative effects and enhancing the robustness of the analysis.
@@
To conclude this section, let’s estimate the feasibility of observing the dead-cone effect with ECF and jet angularities at future $e^{+}\,e^{-}$ colliders. In Fig. 3, we compare our resummed predictions with MC simulations at $\sqrt{s}=2$ TeV center-of-mass collision energy, which corresponds to the ILC setup. The increase in collision energy from 91 GeV to 2 TeV implies that the ratios $\Sigma^{b}/\Sigma^{q}$ in the MC results start to deviate from unity at much smaller values, approximately at $v\sim 0.1$. Similarly, the dead-cone driven changes become visible if $v\lesssim 0.01$, which is in agreement with the leading-log estimate of the dead-cone boundary. These small thresholds pose a significant challenge for future ILC measurements, as detecting the dead cone effect with jets seeded by $b$-quarks would require exceptional detector resolution. A potential solution to this problem could be studying multi-jet events instead of di-jet events, where the average energy per $b$-jet is smaller and hence increases the dead-cone threshold value. Additionally, one could also study the dead cone effect for top quarks, although it’s not yet clear if its extremely short lifetime would allow to produce enough QCD emissions to observe the suppression of collinear radiation. For now, we leave the study of multi-jet ILC events and top production to future work.